OpenStudy (anonymous):

PLEASE HELP Let C(q) represent the total cost of producing q items. Then C′(q) gives the marginal cost in dollars per item. Suppose a company determines that C(13)=2800 and C′(13)=73. Estimate the total cost of producing the following number of items. Your answers should be whole numbers. a) 14 items b) 12 items

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@freckles

ganeshie8 (ganeshie8):

you may use linear approximation

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

Remember, "derivative of a function at a point" gives the "slope of tangent line" at that point.

OpenStudy (anonymous):

f(b + h) ≈ f( b ) + f’( b )∙( h )

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Excellent! so do we get : C(14) = C(13 + 1) = C(13) + C'(13)*(1) ?

OpenStudy (anonymous):

i think so @ganeshie8

ganeshie8 (ganeshie8):

C(13)=2800 and C′(13)=73. plug them in

OpenStudy (anonymous):

ok. is it 2,873?

ganeshie8 (ganeshie8):

Looks good!

OpenStudy (anonymous):

so how would be do b?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@dan815

Nnesha (nnesha):

same steps for b.

OpenStudy (anonymous):

can u help me? @Nnesha